OpenStudy (anonymous):

Which fractions are equivalent to negative 15/20 ? Choose all the answers that are correct A. negative 2/3 B. 9/-12 c. -12/16 D. -25/30

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Well, what is -15/20?

OpenStudy (anonymous):

(divide)

OpenStudy (anonymous):

Oh and \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\)

OpenStudy (anonymous):

Well thanks :) I got -0.75?

OpenStudy (anonymous):

I'm not understanding

OpenStudy (anonymous):

Is it only C?

OpenStudy (kkutie7):

\[-\frac{15}{20}=-.75\] \[-\frac{2}{3}=?\] \[\frac{9}{-12}=?\] \[-\frac{12}{16}=?\] \[-\frac{25}{30}=?\] doing this should help you find the answer